Submarine X23 has parametric equation of 2+t= x(t) & 4-3t= y(t) and Submarine Y18 has parametric equation of 11-t= x(t) & 3+at= y(t). how can i find the time they intersect if Submarine X23 left at 2.17 PM ? and Submarine Y18 leaves 2 min. later ?
Answer by Alan3354(69443)

What is the at ?
11-t= x(t) & 3+at= y(t)